During recess, two out of every three kids were playing soccer. The rest played basketball. If sixty-three kids were at recess, then how many kids were playing basketball?
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


2/3 of the kids played soccer.
The rest (aka 1/3 of the kids) played basketball.


1/3 of 63 = (1/3)*63 = 63/3 = 21 kids were playing basketball.

Answer by ikleyn(52782)

.

Since the problem says "two of every three kids", determine the number of triples.


For it, divide 63 by 3, getting 21 triples.


From each triple, exactly one kid plays basketball.


So, the  ANSWER  is :  21 kids play basketball.